The width of a rectangle is 5 feet less than its length. If the perimeter is 22 feet, what is the width?

Answer by checkley79(3341)

P=2L+2W
W=L-5
22=2L+2(L-5)
22=2L+2L-10
4L=22+10
4L32
L=32/4
L=8 ANS. FOR THE LWNGTH.
W=8-5=3 ANS. FOR THE WIDTH.
PROOF:
22=-2*8+2*3
22=16+6
22=22